4. A conductor of length šæ moves at a steady speed vat right angles to a uniform magnetic fieldĀ 

of flux density šµ. Show that the e.m.f. šø across the ends of the conductor is given by theĀ 

equation šø = šµš‘™š‘£.Ā 


1
Expert's answer
2021-03-30T19:46:38-0400

Consider the area enclosed by the moving rod, rails and resistor. B is perpendicular to this area, and the area is increasing as the rod moves. Thus the magnetic flux enclosed by the rails, rod and resistor is increasing. When flux changes, an EMF is induced according to Faradayā€™s law of induction.

To find the magnitude of EMF induced along the moving rod, we use Faradayā€™s law of induction without the sign:


"EMF=N\frac{\Delta \Phi}{\Delta t}"

In this equation, N=1 and the flux Ī¦=BAcosĪø. We have Īø=0Āŗ and cosĪø=1, since B is perpendicular to A. Now Ī”Ī¦=Ī”(BA)=BĪ”A, since B is uniform. Note that the area swept out by the rod is Ī”A=ā„“x. Entering these quantities into the expression for EMF yields:


"EMF=\frac{B\Delta A}{\Delta t}=\frac{Bl\Delta x}{\Delta t}=Blv"